OpenStudy (anonymous):

In solving the equation (x + 3)(x + 1) = 48, Eric stated that the solution would be x + 3 = 48 => x = 45 or (x + 1) = 48 => x = 47 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning. Will someone please help me

OpenStudy (saifoo.khan):

Discussion problems suck.

OpenStudy (anonymous):

i know... i hate em... its my final exam and im trying to get through been doing it all day

OpenStudy (saifoo.khan):

we have to use the FOIL method to multiply them on the left side of the equation. (x + 3)(x + 1) = 48 \[x^2+4x+3=48 \]\[x^2+4x−45\] When we are going to solve it further by factorization, we will solve it like: \[x^2 + 4x -45 = 0\]\[x^2 + 9x -5x -45 = 0 \] Taking commons: x(x+9) -5(x+9) =0 Therefore: (x+9)(x-5) So solutions are: x = -9 & x = 5

OpenStudy (anonymous):

I was about to do it, got the same answer :)

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

Did taht worked @87??

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

yes... i got one more will u help me?

OpenStudy (saifoo.khan):

Sure, post it on left. ^.^ Dont forget to hit good answer.